# ESP-MedSAM: Efficient Self-Prompting SAM for Universal Domain-Generalized Medical Image Segmentation

### v4

@@ Abstract
The universality of deep neural networks across different modalities and their generalization capabilities to unseen domains play an essential role in medical image segmentation. The recent Segment Anything Model (SAM) has demonstrated its potential in both settings. However, the huge computational costs, demand for manual annotations as prompts and conflict-prone decoding process of SAM degrade its generalizability and applicability in clinical scenarios. To address these issues, we propose an efficient self-prompting SAM for universal domain-generalized medical image segmentation, named ESP-MedSAM. Specifically, we first devise the Multi-Modal Decoupled Knowledge Distillation (MMDKD) strategy to construct a lightweight semi-parameter sharing image encoder that produces discriminative visual features for diverse modalities. Further, we introduce the Self-Patch Prompt Generator (SPPG) to automatically generate high-quality dense prompt embeddings for guiding segmentation decoding. Finally, we design the Query-Decoupled Modality Decoder (QDMD) that leverages a one-to-one strategy to provide an independent decoding channel for every modality. Extensive experiments indicate that ESP-MedSAM outperforms state-of-the-arts in diverse medical imaging segmentation tasks, displaying superior modality universality and generalization capabilities. Especially, ESP-MedSAM uses only 4.5% parameters compared to SAM-H. The source code is available at https://github.com/xq141839/ESP-MedSAM.
{IEEEkeywords}
@@
1 Introduction
@@
Medical imaging has made great strides in the last decades, spawning a variety of modalities, such as histopathology imaging, dermoscopy imaging, X-ray imaging, fundus imaging, colonoscopy imaging and ultrasound imaging. They play an important role in determining disease types and grading [1]. Traditionally, medical images are analyzed by pathologists, which is time-consuming and occupies substantial healthcare resources. In this challenging background, computer-aided diagnosis is expected to accelerate evaluation time and improve diagnostic efficiency, where pixel-level segmentation of target regions is a key step for quantitative and qualitative assessment, presenting valuable information [2].
Convolutional Neural Network (CNN) based U-shape architectures investigate the correlation between the low-level and high-level semantic information in mask prediction [3, 4, 5, 6]. Although these methods demonstrate the accurate generation of segmentation masks within known domains, they are difficult to be generalized to unseen domains. To address Domain Generalization (DG) challenges, existing studies utilize multi-task learning, data augmentation and domain synthesis to improve the diversity of model feature representations [7, 8, 9, 10, 11]. However, these task-specific models require to be retrained from scratch when facing different modalities due to their limited model capacity. This raises significant challenges in establishing a segmentation model with superior generalization and universality across various medical imaging modalities.
Vision Transformer (ViT) [12] leverages the self-attention mechanism to capture long-range dependencies and provides a larger model capacity to overcome the constraints of inductive bias under the supervision of big data. Especially, the recent appearance of Segment Anything Model (SAM) [13] has made a significant breakthrough in the field of image segmentation. The superior generalization of SAM in natural images has been sufficiently validated, demonstrating its transferability in a wide range of scenarios through interactive prompts. Due to its pre-trained large-scale image encoder, current studies [14, 15] have illustrated the potential of SAM for universal domain-generalized universal medical image segmentation. Despite these advantages, adapting SAM to clinical scenarios faces three significant obstacles. Firstly, the standard SAM [13] contains a large number of parameters. Note that the SAM-H contains 636M parameters. The huge computational costs limit the applicability of SAM in real-world scenarios. Although existing methods based on Parameter-Efficient Fine-Tuning (PEFT) techniques [16, 17] enable the reduction of learnable parameters during the fine-tuning phase, the size of the entire model is not decreased, but even increased.
Furthermore, SAM [13] mainly depends on manual annotations (e.g., points, boxes) as prompts to guide the segmentation mask generation. Although the SAM based on the bounding box prompt mode can produce relatively precise masks, it requires users to roughly provide the location of diseases in images, which can only be completed by medical experts, defeating the purpose of healthcare AI. Particularly, for the most effortless mode (i.e., single positive point) and simple automatic prompt generation algorithm, recent studies [14, 15] have indicated that they are difficult to perform acceptable results in medical applications due to insufficient or incorrect prompt information. Moreover, the original SAM [13] leverages the (modality-agnostic mask queries to predict potential segmentation masks for all categories. However, in the field of medical imaging, every vision modality has inherent heterogeneity. This many-to-many strategy is difficult to handle the mutual knowledge interference of different modalities, resulting in the degradation of model generalization.
In this paper, we propose an efficient self-prompting SAM framework for universal domain-generalized medical image segmentation, named ESP-MedSAM. Specifically, we first introduce the Multi-Modal Decoupled Knowledge Distillation (MMDKD) method to build a lightweight semi-parameter sharing image encoder that employs the modality-specific knowledge to drive the modality aggregator producing discriminative visual features across diverse modalities. Then, we devise the Self-Patch Prompt Generator (SPPG) to automatically produce a set of high-quality patch prompts assisting the prediction of segmentation masks. Moreover, we introduce the Query-Decoupled Modality Decoder (QDMD), applying the one-to-one strategy to personalise the decoding process for each medical modality, preventing the different modalities from interfering with each other.
The contributions of this work are summarized as follows.
  • We propose the MMDKD strategy to distil the modality-specific and modality-common knowledge from the foundation model to a modality controller and modality aggregator, respectively. Both constitute a lightweight semi-parameter sharing image encoder, leading to the generation of distinguished features for different medical modalities.
  • We devise the SPPG to produce high-quality patch prompts automatically without the demand for manual annotations. These prompts are used to guide the prediction of segmentation masks.
  • We introduce the QDMD for the segmentation decoding. It leverages the one-to-one strategy to provide a private segmentation workflow to each modality, preventing the different modalities from interfering with each other.
  • We take the distilled lightweight semi-parameter sharing image encoder, SPPG and QDMD to build our ESP-MedSAM with remarkable generalization-efficiency trade-offs. We conduct extensive experiments on diverse medical imaging modalities, proving that our ESP-MedSAM outperforms state-of-the-arts in universal domain-generalized medical image segmentation.
Figure 1: The end-to-end architecture of our ESP-MedSAM for universal medical image segmentation. (a) Multi-Modal Decoupled Knowledge Distillation. (b) Self-Patch Prompt Generator. (c) Query-Decoupled Modality Decoder.
2 Related Work
2.1 Generalized Medical Image Segmentation
The generalization capabilities of deep neural networks have received significant attention in medical image segmentation. The original UNet [3] reveals great single-domain adaptation but is difficult to generalize unseen domains. Previous studies [4, 18, 7] mainly adopt multi-scale feature fusion to improve the feature representation power of models. In addition, Cheng et al. [8] utilized causality-inspired data augmentation to extend the distribution of the single-source domain during training. Xu et al. [9] proposed an Adversarial Domain Synthesizer (ADS) to synthesize the new domains from the memorized source domain information. As these task-specific methods have limited model capacity, they need to be trained from scratch for each modality. The recent Segment Anything Model (SAM) [13] took advantages of its large-scale image encoder and interactive prompts to achieve outstanding zero-shot generalization in natural image segmentation. In medical image segmentation, MedSAM [15] and SAMMI [14] collected more than 1M public medical images to fully fine-tune SAM with box and point prompts for domain generalized universal medical image segmentation. However, such methods rapidly increase data and computation costs, which are expensive and impractical in clinical scenarios. To mitigate the reliance on data size and computational resources during transfer learning, parameter-efficient fine-tuning techniques have been introduced in SAM. Specifically, Adapter has been widely used to integrate into the image encoder of SAM for the refinement of feature representation in medical imaging [17, 16]. SAMed [19] concatenated Low-Rank Adaptation (LoRA) with self-attention layers of SAM to optimize the feature extraction. Zhong et al. [20] proposed Conv-LoRA that incorporated multiple parallel convolutional experts with LoRA to sample different feature scales. While beneficial, these architectures are still based on the huge ViT encoder, which is computationally expensive. In contrast, our approach overcomes this aforementioned challenge and illustrates superior generalization-efficiency trade-offs across a variety of medical imaging modalities.
Furthermore, SAM [13] relying on laborious manual annotations as segmentation prompts seriously reduces its applicability in clinical scenarios. Although it provides a simple sliding window algorithm to automatically generate box and centre point prompts from inputs, recent studies [14, 21] have indicated that this approach fails to perform satisfactory results in medical image segmentation tasks. To minimize the need for manual annotations, SPPNet [22] computed neighbouring points around the centroid as extra prompts for nuclei image segmentation. Moreover, many studies used traditional segmentation networks (e.g., UNet [3]) to produce low-resolution masks as prompts [23, 24, 25]. However, such pixel-level small segmentors may generate more error prompts when facing heterogeneous modalities due to their limited model capacity. On the contrary, our ESP-MedSAM framework automatically generates a set of high-quality patch prompts from its own image embeddings for guiding segmentation decoding, so it eliminates the demand for labour-intensive manual annotations.
2.2 Knowledge Distillation
Knowledge Distillation (KD) [26] is a method of compressing the size of foundation models, transferring the knowledge from the teacher model to the student model. Generally, hard labels (e.g., category) and soft labels (e.g., probability) are used to supervise the learning of a student model from a teacher model. Yang et al. [27] decoupled the distillation into two stages: representation learning and classification. Subsequently, decoupled knowledge distillation [28] was introduced to divide the traditional KD loss into two parts: target class and non-target class knowledge distillation, which enhances the efficiency of knowledge transfer between the teacher and student model. In addition, it has been proven that soft labels are preferable in KD-based medical image segmentation [29, 30, 31, 32]. Recently, transferring the knowledge of SAM to a small model has become a hot research topic. MobileSAM [33] retained the prompt encoder and mask decoder of the standard SAM [13], adopting feature distillation between its image encoder and the TinyViT [34]. EdgeSAM [35] involved both the prompt encoder and mask decoder in the distillation process to capture the full knowledge embodied in SAM. EfficientSAM [36] utilised masked image pretraining method and reconstruction loss to transfer the knowledge from the image encoder of SAM to a lightweight encoder. However, these feature-coupled distillation methods are challenging to harmonise the feature representation across diverse modalities with inherent heterogeneity, resulting in the degradation of the model generalization capability. On the contrary, our method adopts a feature-level decoupled distillation strategy. The distilled image encoder enables the generation of discriminative features for different medical modalities.
3 Methodology
3.1 Overview of ESP-MedSAM
In this study, we denote $\mathcal{S}=\{\mathcal{S}^{1},\mathcal{S}^{2},\cdots,\mathcal{S}^{K}\}$ as the set of $K$ highly heterogeneous medical vision modalities involved in source domains. Each domain includes image and segmentation mask pairs of $\mathcal{S}^{k}=\{(x_{i},y_{i})\}_{i=1}^{N}$. Our goal is to train a universal segmentation model $\mathcal{F}_{\theta}:x\rightarrow y$ on source domains, which not only performs well across different modalities but also can be directly generalized to an unseen target domain $\mathcal{T}^{k}$ of each modality.
As presented in Fig. 1, we illustrate our ESP-MedSAM architecture for domain-generalized medical image segmentation across various modalities. Given medical images from the $k$-th modality, we first utilize a lightweight semi-parameter sharing image encoder to produce a set of discriminative image embeddings. This module contains two subcomponents: modality controller and modality aggregator. Both are distilled from the large ViT-based teacher model using the MMDKD strategy. Then, these embeddings are delivered to the SPPG module, which is automatically converted to a set of high-quality patch prompts to assist segmentation decoding. Finally, QDMD leverages the independent query of the corresponding modality to customize the decoding process for individual medical modalities.
3.2 Multi-Modal Decoupled Knowledge Distillation
SAM [13] mainly relies on a large-capacity ViT to provide generalized feature representations but is challenged by the huge computation costs, which limit its applications in real-world scenarios. Recent studies [35, 36, 37, 33, 38, 39] mainly aim to transfer the knowledge of the huge image encoder into a lightweight architecture. However, different modalities of medical imaging face inherent heterogeneity. With the reduction of model capacity, such full-parameter sharing image encoders are difficult to harmonise the feature representation of the model across diverse modalities, thereby degrading the model generalization. To address the aforementioned challenge, we propose the MMDKD to transfer the knowledge of foundation models to a lightweight semi-parameter sharing image encoder that enables generating discriminative features for different modalities, as presented in Fig. 1a.
3.2.1 Stucture of Multi-Modal Teacher Model
To conduct the distillation process, we first construct a multi-modal teacher model that involves the natural image encoder $\mathcal{F}_{\rm{ViT\text{-}SAM}}$ of SAM [13] and a medical image encoder $\mathcal{F}_{\rm{ViT\text{-}Med}}$. Specifically, we freeze the weight of the image encoder in SAM and additionally create a set of learnable tokens $\{\theta_{\rm key}^{\rm tr},\theta_{\rm value}^{\rm tr}\}$, where $\theta^{\rm tr}\in\mathbb{R}^{t\times d}$, $t$ and $d$ refer to the number and dimension of tokens, respectively. They are concatenated with the key and value channels of the multi-head attention layer to learn new patterns in medical images, updating attention maps. Further, we parallelize learnable multi-layer perceptron $\Phi_{\rm tr}$ to the Feed Forward Network (FFN) in ViT for storing diverse modality information. We take this $\mathcal{F}_{\rm{ViT\text{-}Med}}$, SPPG and QDMD to implement fine-tuning on our defined medical source domains $\mathcal{S}$. The teacher model is supervised by a combination of focal loss and dice loss:
$L_{\rm Mask}=\lambda L_{\rm dice}+(1-\lambda)L_{\rm focal},$ (1)
where $\lambda$ represents the coefficient to adjust the weight of two loss functions. As a result, $\mathcal{F}_{\rm{ViT\text{-}SAM}}$ and $\mathcal{F}_{\rm{ViT\text{-}Med}}$ enable providing modality-common knowledge and modality-specific knowledge, respectively.
3.2.2 Decoupled Knowledge Distillation
The primary challenge in performing KD from medical foundation models to lightweight frameworks is avoiding the mutual knowledge interference of different modalities with inherent heterogeneity. Intuitively, the quality of segmentation masks depends on the feature extracted from the image encoder. Drawing inspiration from the divide-and-conquer algorithm, our proposed MMDKD method decouples the KD process into two sub-tasks: modality-common feature distillation and modality-specific feature distillation. During modality-common feature distillation, We aim to transfer the knowledge from $\mathcal{F}_{\rm{ViT\text{-}SAM}}$ to a modality aggregator $\mathcal{F}_{\rm{MA}}$. It is composed of the weights from the small ViT (e.g., TinyViT [34]). Moreover, we distil the specific knowledge of each medical modality from $\mathcal{F}_{\rm{ViT\text{-}Med}}$ to the corresponding parameter bags $\{w^{k}_{\rm key},w^{k}_{\rm value},\Phi^{k}_{\rm expert}\}$, where $w^{k}_{\rm key},w^{k}_{\rm value}\in\mathbb{R}^{z\times d}$ are trainable tokens that respectively store the knowledge of key and value branches in $\mathcal{F}_{\rm ViT-Med}$, and $\Phi^{k}_{\rm expert}$ is MLP, with $k\in 1,2,\cdots,K$ and $z$ being the number of tokens. All parameter bags constitute a modality controller $\mathcal{F}_{\rm{MC}}$. The loss function of optimising these two modules is defined by:
$L_{\rm MMDKD}$ $=\big{|}\big{|}\mathcal{F}_{\rm{ViT\text{-}SAM}}(\textbf{{x}}_{\rm nat})-\mathcal{F}_{\rm{MA}}(\textbf{{x}}_{\rm nat})\big{|}\big{|}^{2}_{2}$ (2)
$+\sum^{K}_{k=1}\big{|}\big{|}\mathcal{F}_{\rm{ViT\text{-}Med}}(\textbf{{x}}^{k})-\mathcal{F}_{\rm{MC}}(\textbf{{x}}^{k})\big{|}\big{|}^{2}_{2},$
where $\textbf{{x}}_{\rm nat}$ stands for the $1\%$ natural images of the SA-1B dataset [13] and $\textbf{{x}}^{k}$ is the medical image collected from source domains $\mathcal{S}^{k}$. The MMDKD strategy provides a private distillation channel for each medical imaging modality, preventing different modalities from interfering with each other.
3.2.3 Lightweight Semi-Parameter Sharing Image Encoder
To seamlessly harmonize feature representations across diverse modalities, we take the distilled modality controller and modality aggregator to build a lightweight semi-parameter sharing image encoder. Given a medical image $x^{k}_{i}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$ and $C$ are the height, width and number of channels of the image, our modality controller first send $k$-th parameter bag to $\mathcal{F}_{\rm{MA}}$. Specifically, $\{w^{k}_{\rm key},w^{k}_{\rm value}\}$ concatenates with $\mathcal{K}$ and $\mathcal{V}$ branches of the modality aggregator to adapt the attention computation $\mathcal{A}$ of the specific modality:
$\mathcal{A}=\mathrm{softmax}(\frac{\mathcal{Q}(\mu^{k}_{i})\cdot{(\mathcal{K}(\mu^{k}_{i})\smallfrown w^{k}_{\rm key})}^{T}}{\sqrt{d}})\cdot(\mathcal{V}(\mu^{k}_{i})\smallfrown w^{k}_{\rm value}),$ (3)
[…]
$\textit{h}=\mu^{k}_{i}+\mathcal{A}+\Phi_{\rm agg}(\mathcal{F}_{\rm LN}(\mathcal{A}))+\Phi^{k}_{\rm expert}(\mathcal{F}_{\rm LN}(\mathcal{A})),$ (4)
where h is a set of image embeddings and $\mathcal{F}_{\rm LN}$ is LayerNorm. Overall, our proposed MMDKD method efficiently decreases the computational costs of the image encoder compared to SAM while retaining the ability to produce generalizable feature maps.
Figure 2: The illustration of Self-Patch Prompt Generator.
3.3 Self-Patch Prompt Generator

### v5

# De-LightSAM: Modality-Decoupled Lightweight SAM for Generalizable Medical Segmentation

@@ Abstract
The universality of deep neural networks across different modalities and their generalization capabilities to unseen domains play an essential role in medical image segmentation. The recent segment anything model (SAM) has demonstrated strong adaptability across diverse natural scenarios. However, the huge computational costs, demand for manual annotations as prompts and conflict-prone decoding process of SAM degrade its generalization capabilities in medical scenarios. To address these limitations, we propose a modality-decoupled lightweight SAM for domain-generalized medical image segmentation, named De-LightSAM. Specifically, we first devise a lightweight domain-controllable image encoder (DC-Encoder) that produces discriminative visual features for diverse modalities. Further, we introduce the self-patch prompt generator (SP-Generator) to automatically generate high-quality dense prompt embeddings for guiding segmentation decoding. Finally, we design the query-decoupled modality decoder (QM-Decoder) that leverages a one-to-one strategy to provide an independent decoding channel for every modality, preventing mutual knowledge interference of different modalities. Moreover, we design a multi-modal decoupled knowledge distillation (MDKD) strategy to leverage robust common knowledge to complement domain-specific medical feature representations. Extensive experiments indicate that De-LightSAM outperforms state-of-the-arts in diverse medical imaging segmentation tasks, displaying superior modality universality and generalization capabilities. Especially, De-LightSAM uses only 2.0% parameters compared to SAM-H. The source code is available at https://github.com/xq141839/De-LightSAM.
Index Terms:
@@
I Introduction
Medical imaging has made great strides in the last decades, spawning a variety of modalities, such as histopathology imaging, dermoscopy imaging, X-ray imaging, fundus imaging, colonoscopy imaging and ultrasound imaging. They play an important role in determining disease types and grading [1, 2, 3]. Traditionally, medical images are analyzed by medical experts, which is time-consuming and occupies substantial healthcare resources. In this challenging background, computer-aided diagnosis is expected to accelerate evaluation time and improve diagnostic efficiency, where pixel-level segmentation of target regions is a key step for quantitative and qualitative assessment, presenting valuable information [4].
Figure 1: Comparison of our De-LightSAM and existing medical SAM works. (a) Small Unet-like models for automatic prompt generation. (b) Distilled image encoders for lower computation cost. (c) Our De-LightSAM achieves efficient computation cost and auto-prompting generation simultaneously.
Convolutional neural network (CNN) based U-shape architectures investigate the correlation between the low-level and high-level semantic information in mask prediction [5, 6, 7, 8, 9, 10, 11]. Although these methods demonstrate the accurate generation of segmentation masks within known domains, they are difficult to generalize to unseen domains. To address domain generalization (DG) challenges, existing studies utilize multi-task learning, data augmentation and domain synthesis to improve the diversity of model feature representations [12, 13, 14]. However, these task-specific models require to be retrained from scratch when facing different modalities due to their limited model capacity. This raises significant challenges in establishing a segmentation model with superior generalization and universality in various medical imaging modalities.
Vision transformer (ViT) [15] leverages the self-attention mechanism to capture long-range dependencies and provides a larger model capacity to overcome the constraints of inductive bias under the supervision of big data. Especially, the recent appearance of segment anything model (SAM) [16] has made a significant breakthrough in the field of image segmentation. The superior generalization of SAM in natural images has been sufficiently validated, demonstrating its transferability in a wide range of scenarios through interactive prompts. Existing studies [17, 18] have illustrated the potential of SAM for universal medical image segmentation. Despite these advantages, adapting SAM to clinical scenarios faces three significant obstacles. Firstly, SAM [16] mainly depends on manual annotations (e.g., points, boxes) as prompts to guide the segmentation mask generation, leading to labor-intensive operations. Particularly, for the most effortless mode (i.e., single positive point) and simple automatic prompt generation algorithm (i.e., sliding window), recent studies [17, 18] have indicated that they are difficult to perform acceptable results in medical applications due to insufficient or incorrect prompt information. Although current studies [19, 20, 21] include classical small Unet or object detection models, as shown in Fig. 1(a), to automatically generate different prompts, they involve significant additional computation cost.
Furthermore, the standard SAM [16] contains a large number of parameters. In particular, the image encoder of SAM-H contains 636M parameters. The huge computational costs limit the applicability of SAM in real-world scenarios. Existing methods based on parameter-efficient fine-tuning techniques [4] enable the reduction of learnable parameters during the fine-tuning phase. In addition, to reduce inference cost, many studies [22, 23, 24] leverage knowledge distillation strategies to transfer the knowledge of SAM to a single small image encoder, as illustrated in Fig. 1(b). Meanwhile, they leverage a modality-agnostic decoding query to predict potential segmentation masks for all categories. However, in the field of medical imaging, every vision modality has inherent heterogeneity. This one-to-many strategy is difficult to handle the mutual knowledge interference of different modalities, resulting in the degradation of model generalization capabilities.
To address these limitations, we propose De-LightSAM, a modality-decoupled lightweight SAM framework specifically designed for domain-generalized medical image segmentation, as shown in Fig. 1(b) Specifically, we first develop a lightweight domain-controllable image encoder (DC-Encoder) that efficiently captures discriminative visual features across diverse medical modalities while significantly reducing computational overhead. Second, we introduce a self-patch prompt generator (SP-Generator) that eliminates the need for manual annotations by automatically generating high-quality dense prompt embeddings, enabling fully automated segmentation. Third, we design a query-decoupled modality decoder (QM-Decoder) that employs a one-to-one decoding strategy, providing independent channels for each modality to prevent cross-modal knowledge interference and enhance segmentation accuracy. Additionally, we propose a multi-modal decoupled knowledge distillation (MDKD) strategy that effectively transfers both common and domain-specific knowledge from foundation models to our lightweight architecture. Through extensive experiments across multiple medical imaging datasets, we demonstrate that De-LightSAM achieves superior performance compared to state-of-the-art methods while using only 2.0% of SAM-H’s parameters, showcasing exceptional computation efficiency and cross-domain generalization capabilities.
The contributions of this work are summarized as follows.
  • We propose De-LightSAM as an end-to-end framework that addresses real-world medical imaging challenges, including computational constraints, and multi-modal compatibility. De-LightSAM offers a practical pathway in clinical environments with limited computational resources while maintaining high segmentation accuracy across diverse medical imaging scenarios.
  • We devise a DC-Encoder integrated with the MDKD strategy. This approach efficiently extracts discriminative visual features across diverse medical modalities while transferring both common and domain-specific knowledge from foundation models, achieving the significant reduction of computation cost.
  • We introduce an integrated framework combining an SP-Generator for automatic high-quality prompt generation and a QM-Decoder with a one-to-one decoding strategy. This eliminates manual annotation dependency while providing independent processing channels for each modality to prevent cross-modal interference.
  • We conduct extensive experiments across multiple medical imaging modalities and datasets to systematically evaluate domain generalization capabilities. Our comprehensive analysis demonstrates superior generalization capabilities compared to existing state-of-the-art methods.
Figure 2: The overview of our De-LightSAM framework for domain-generalized medical image segmentation, consisting of DC-Encoder, SP-Generator and QM-Decoder. For ease of understanding, we elaborate on the case of De-LightSAM in skin lesion segmentation. Our De-LightSAM fully exploits a modality-decoupled paradigm to enhance generalization capabilities.
II Related Work
II-A Generalized Medical Image Segmentation
The generalization capabilities of deep neural networks have received significant attention in medical image segmentation. The original UNet [5] reveals great single-domain adaptation but is difficult to generalize unseen domains. Previous studies [6, 25, 12, 26, 27] mainly adopt multi-scale feature fusion to improve the feature representation power of models. In addition, Cheng et al. [13] utilized causality-inspired data augmentation to extend the distribution of the single-source domain during training. Xu et al. [14] proposed an Adversarial Domain Synthesizer (ADS) to synthesize the new domains from the memorized source domain information. As these task-specific methods have limited model capacity, they need to be trained from scratch for each modality. The recent Segment Anything Model (SAM) [16] took advantages of its large-scale image encoder and interactive prompts to achieve outstanding zero-shot generalization in natural image segmentation. In medical image segmentation, MedSAM [18] and SAMMI [17] collected more than 1M public medical images to fully fine-tune SAM with box and point prompts for universal medical image segmentation. However, such methods rapidly increase data and computation costs, which are expensive and impractical in clinical scenarios. To mitigate the reliance on data size and computational resources during transfer learning, parameter-efficient fine-tuning techniques have been introduced in SAM. Specifically, the Adapter has been widely used to integrate into the image encoder of SAM for the refinement of feature representation in medical imaging [4]. SAMed [28] concatenated Low-Rank Adaptation (LoRA) with self-attention layers of SAM to optimize the feature extraction. While beneficial, these architectures are still based on the huge ViT encoder, which is computationally expensive. In contrast, our approach overcomes this aforementioned challenge and illustrates superior generalization-efficiency trade-offs across a variety of medical imaging modalities.
Furthermore, SAM [16] relying on laborious manual annotations as segmentation prompts seriously reduces its applicability in clinical scenarios. Although it provides a simple sliding window algorithm to automatically generate box and center point prompts from inputs, recent studies [17, 29] have indicated that this approach significantly degrades inference speed and fails to perform satisfactory results in medical image segmentation tasks. To minimize the need for manual annotations, many studies used traditional segmentation networks (e.g., UNet [5]) to produce low-resolution masks as prompts [20, 30]. However, such pixel-level small segmentors may generate more error prompts when facing heterogeneous modalities due to their limited model capacity. On the contrary, our De-LightSAM framework automatically generates a set of high-quality patch prompts from its own image embeddings for guiding segmentation decoding, so it eliminates the demand for labour-intensive manual annotations.
II-B Knowledge Distillation
Knowledge distillation (KD) [31] is a classical method of compressing the size of foundation models, transferring the knowledge from the teacher model to the student model. Generally, hard labels (e.g., category) and soft labels (e.g., probability) are used to supervise the learning of a student model from a teacher model. Yang et al. [32] decoupled the distillation into two stages: representation learning and classification. Subsequently, decoupled knowledge distillation [33] was introduced to divide the traditional KD loss into two parts: target class and non-target class knowledge distillation, which enhances the efficiency of knowledge transfer between the teacher and student model. In addition, it has been proven that soft labels are preferable in KD-based medical image segmentation [34, 35]. Recently, transferring the knowledge of SAM to a small model has become a hot research topic. MobileSAM [22] retained the prompt encoder and mask decoder of the standard SAM [16], adopting feature distillation between its image encoder and the TinyViT [36]. EdgeSAM [37] involved both the prompt encoder and mask decoder in the distillation process to capture the full knowledge embodied in SAM. EfficientSAM [23] utilized masked image pretraining method and reconstruction loss to transfer the knowledge from the image encoder of SAM to a lightweight encoder. However, these feature-coupled distillation methods are challenging to harmonize the feature representation across diverse modalities with inherent heterogeneity, resulting in the degradation of the model generalization capability. On the contrary, our method adopts a feature-level decoupled distillation strategy. The distilled image encoder enables the generation of discriminative features for different medical modalities.
Figure 3: The illustration of our MDKD strategy, including two sub-tasks: common feature distillation and domain-specific feature distillation. We leverage robust common knowledge to complement domain-specific medical feature representations.
III Methodology
III-A Overview of De-LightSAM
In this study, we denote $\mathcal{S}=\{\mathcal{S}^{1},\mathcal{S}^{2},\cdots,\mathcal{S}^{K}\}$ as the set of $K$ highly heterogeneous medical vision modalities involved in source domains. Each domain includes image and segmentation mask pairs of $\mathcal{S}^{k}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $N$ represents the number of pairs within the domain. Our goal is to train a universal segmentation model $\mathcal{F}_{\theta}:x\rightarrow y$ on source domains, which not only performs well across different modalities but also can be directly generalized to an unseen target domain $\mathcal{T}^{k}$ of each modality.
As illustrated in Fig. 2, De-LightSAM presents a comprehensive framework for domain-generalized medical image segmentation that operates efficiently across diverse imaging modalities. Given input medical images from the $k$-th modality, our framework processes segmentation through three core components. Firstly, the DC-Encoder extracts discriminative visual representations while maintaining modality-specific characteristics. This lightweight encoder incorporates a modality controller for domain-aware feature learning, optimized through our proposed MDKD strategy from pre-trained foundation models. Then, the SP-Generator automatically transforms the encoded features into high-quality dense prompt embeddings, eliminating the dependency on manual annotations while ensuring comprehensive spatial coverage for accurate segmentation guidance. Finally, the QDMD employs modality-specific decoding channels, where each modality utilizes its dedicated query mechanism to perform independent segmentation inference, thereby preventing cross-modal interference and enhancing modality-specific performance. This end-to-end architecture enables De-LightSAM to achieve superior generalization across unseen medical domains while maintaining computational efficiency.
III-B Domain-Controllable Image Encoder
SAM [16] mainly relies on a large-capacity ViT to provide generalized feature representations but is challenged by the huge computation costs, which limit its applications in real-world scenarios. Recent studies [38, 24, 39] mainly aim to transfer the knowledge of the huge image encoder into an independent lightweight architecture. However, different modalities of medical imaging face inherent heterogeneity. Such full-knowledge sharing image encoders are difficult to harmonize the feature representation of the model across diverse modalities, degrading the generalization capabilities. To address this challenge, we propose a DC-Encoder that generates discriminative features for different modalities while maintaining computational efficiency. Given a set of patch embeddings $x^{k}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$ and $C_{1}$ are the height, width and number of channels, our DC-Encoder first employs two MBConvs [40] blocks to learn low-level representation efficiently. Specifically, each MBConv block applies a $1\times 1$ convolution for channel expansion. A $3\times 3$ depthwise convolution is followed by a $1\times 1$ projection convolution for effective channel communication. Then, MBConv blocks combine with $L_{1}$ standard ViT layers to capture long-range dependencies. In particular, we create a modality controller $\mathcal{F}_{\rm{MC}}=\{w^{k}_{\rm bias},w^{k}_{\rm up},w^{k}_{\rm down}\}_{k=1}^{K}$, where $w^{k}_{\rm bias}$ and $w^{k}_{\rm MLP}$ represent learnable tokens and MLP layers, respectively. Every $w^{k}_{\rm bias}$ concatenates with the query $\mathcal{Q}$ branche of the self-attention layer to adapt the attention computation $\mathcal{A}$ of the specific modality:
$\mathcal{A}=\mathrm{softmax}(\frac{\mathcal{Q}(x^{k}\smallfrown w^{k}_{\rm bias})\cdot{\mathcal{K}(x^{k})}^{T}}{\sqrt{d}})\cdot\mathcal{V}(x^{k}),$ (1)
@@ III-B Domain-Controllable Image Encoder
$\textit{h}=x^{k}+\mathcal{A}+\mathcal{F}_{\rm FFN}(\mathcal{F}_{\rm LN}(\mathcal{A}))+f(w^{k}_{\rm up}\mathcal{F}_{\rm LN}(\mathcal{A}))w^{k}_{\rm down},$ (2)
where $f(\cdot)$ is the activation layer (e.g., GELU) and $\mathcal{F}_{\rm LN}(\cdot)$ is the LayerNorm operation. In addition, the embedding dimension width $C_{1}$ and network depth $L_{1}$ are critical factors affecting computational cost. Previous research [36] has demonstrated that wider dimensions facilitate learning of complex representations, while deeper models may lead to overfitting on medical datasets due to limited fully-annotated labels caused by expensive pixel-level annotation costs. To build a lightweight yet effective architecture, our DC-Encoder adopts a balanced design with moderate embedding dimensions and shallow structure (i.e., $C_{1}=192,L_{1}=10$). This design choice ensures that Encoder can effectively capture domain-specific features while maintaining computational efficiency.
III-C Self-Patch Prompt Generator